OpenStudy (anonymous):

what number must you add to complete the square x^2+12x=-3

OpenStudy (anonymous):

add 36 on both sides to get (x+6)^2=33

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

see basically the logic is to bring it to the form (x+b)^2=x^2+b2+2xb, so for getting this factorisation, see appropriately what to add. observe that the x term has co-efficient 2b and so if u hav a b^2 term u will form the complete square by the formula rite? so u add b^2